OpenStudy (anonymous):

Need help with this.... Solve the following system of equations using the substitution method. x+3y=10 xy= 7

OpenStudy (anonymous):

Isolate one of the variables in one of the equations. I suggest x in the first equation x+3y = 10 subtract 3y from both sides x = 10-3y now substitute this x in the second equation xy = 7 becomes (10-3y) y = 7 expand 10y - 3y^2 = 7 subtract 7 from both sides 10y - 3y^2 - 7 = 0 multiply by -1 3y^2 -10 y + 7 = 0 Factor: (3y-7) (y-1) = 0 y = 7/3 or 1 Plug these into the original equation x = 10-3y x = 3 or x = 7

OpenStudy (baldymcgee6):

x+3y=10 x = 10 -3y SUB INTO NEXT EQUATION (10 - 3y)y = 7 10y - 3y^2 -7 = 0 -3y^2 +10y - 7 = 0 FACTOR

OpenStudy (anonymous):

Thank you! I understand it now.